Morton Companys contribution format income statement for last month is given below: Sales (41,000 units $23 per unit) Variable expenses Contribution margin Fixed expenses Net operating income S 943,000 660,100 282,900 226,320 56,580 The industry in which Morton Company operates is quite sensitive to cyclical movements in the economy. Thus, profits vary considerably from year to year according to general economic conditions. The company has a large amount of unused capacity and is studying ways of improving profits. Required 1. New equipment has come onto the market that would allow Morton Company to automate a portion of its operations. Variable expenses would be reduced by $6.90 per unit. However, fixed expenses would increase to a total of $509,220 each month. Prepare two contribution format income statements, one showing present operations and one showing how operations would appear if the new equipment is purchased. (Round your Per unit answers to 2 decimal places.) Morton Company Contribution Income Statement Present Proposed Amount Per Unit Amount Per Unit 23.0of 1001%! 9.20 13.80 Sales S 943,000 S 23.00 16.10 6.90 660,100 282,900S 226,320 377,200 565,800 $ 509,220 56,580 70 % 40 % riable expenses Contribution margin Fixed expenses 301% 601% t operating income 56,580
Answer #1

Solution 1:

Morton Company
Contribution Income Statement
Particulars Present Proposed
Amount Per unit % Amount Per unit %
Sales $943,000.00 $23.00 100% $943,000.00 $23.00 100%
Variable Cost $660,100.00 $16.10 70% $377,200.00 $9.20 40%
Contribution margin $282,900.00 $6.90 30% $565,800.00 $13.80 60%
Fixed expenses $226,320.00 $509,220.00
Net Operating Income $56,580.00 $56,580.00

Solution 2 a:

Operating leverage = Contribution /Operating income

Present condition = $282900 / $56,580 = 5

Proposed condition = $565,800 / $56,580 = 10

Solution 2b:

Breakeven point in dollar sale = Fixed cost / contribution margin ratio

Present = $226,320 / 30% = $754,400

Proposed = $509,220 / 60% = $848,700

Solution 2c:

Margin of safety sale = Current sales - Breakeven sales

Present = $943,000 - $754,400 = $188,600

Proposed = $943,000 - $848,700 = $94,300

Margin of safety percentage = Margin of safety sales / Current sales

Present = $188,600 / $943,000 = 20%

Proposed = $94,300 / $943,000 = 10%

Solution 3:

Cyclical movement in the economy paramount in mind in deciding to purchase the equipment.

Solution 4:

New net opeating income = $56,580 * 125% = $70,725

New contribution margin = $70,725 + $282,900 = $353,625

New Sales = $943,000 * 150% = $1,414,500

New contribution margin ratio = $353,625 / $1,414,500 = 25%

Breakeven point in dollar sale for new marketing strategy = $282,900 / 25% = $1,131,600